# Surf-D: Generating High-Quality Surfaces of Arbitrary Topologies Using Diffusion Models

### v2

@@ 1 Introduction
Generating a diverse array of 3D shapes using various conditions is an important problem in computer vision and graphics, given its wide applications, such as gaming, filmmaking, and the metaverse. Recent advancements for 3D shape generation [5, 1, 41, 30, 20, 33, 39] aim to make this process more accessible.
The diverse interplay of topology and intricate geometry gives rise to countless 3D shapes in our world. However, existing generation works still suffer from limited topologies and the lack of geometric details. Previous methods present shapes using point clouds [1, 41], voxel representations [58, 72], and signed distance fields (SDFs) [11, 45].
@@
However, point cloud and voxel representations encounter difficulties in faithfully extracting surfaces. Implicit representations, particularly SDFs, excel in describing closed surfaces but face challenges with open surfaces and complex topologies. Despite attempts to enhance them, e.g. 3PSDF [10] and TSDF [61], the representations remain reliant on the underlying definition of SDF. G-Shell [35] is proposed to represent open surfaces but still falters with non-manifold structures and computational constraints.
@@
Moreover, to learn the compact shape latent space efficiently and effectively on diverse surfaces, we employ a curriculum learning strategy [2] to train our model in an easy-to-hard sampling order, i.e. progressively extending the current training dataset during training. This strategy promotes our model to develop a diverse latent space, avoiding entrapment into generating mean shape, i.e. model collapse. As a result, our model is not only much more lightweight and learning-efficient but also feasible to convert high-quality generated implicit shapes to explicit meshes.
@@
    Surf-D achieves superior performance in generating 3D shapes with various topologies while enabling various conditional generation tasks with different modalities.
@@ 2 Related Work
Tremendous efforts have been made for 3D shape generation [5, 1, 41, 4, 20, 34, 7], given its broad applications, employing various shape representations such as voxel grids [70, 68], point clouds [41, 64, 78], and implicit neural fields [48, 17, 12, 24, 40]. Additionally, various methods have been developed to guide the synthesis with text prompt [55, 18, 37, 9, 36], single-view image [33, 39, 32, 31], sketch [71, 76] and sparse control anchors [42, 13]. In the early stages, voxel-based generation methods [70, 68, 51] can synthesize various coarse shapes. However, the voxel representation can not scale to high-resolution generation [77, 76], due to its cubic space complexity. Utilizing point clouds [41, 64, 78], which are compact and readily available, for 3D object representation can achieve arbitrary shape generation with their well-suitability to learn geometry and texture information [46, 71]. However, it requires complex post-processing to faithfully convert the synthesized point clouds into meshes for real-world applications [49].
@@ 2.2 Diffusion Probabilistic Models
Diffusion probabilistic models [23] have demonstrated superior performance in generating samples from a distribution by learning to denoise from a data point gradually [16, 66, 65, 63]. The diffusion models produce high-quality contents including images [33, 16, 47, 66, 53], video [65, 22, 19], and human motion [63, 25, 56]. Previous shape generation works [34, 24] directly train diffusion models on massive grids storing shape representation like SDF, consuming plenty of GPUs and taking days for training. To improve the efficiency, the latent diffusion approach [53] has been explored in 3D shape generation. However, naively applying the latent diffusion model leads to poor generation capability as the vanilla neural field is easy to overfit a single shape. Thus, many efficient generation frameworks are proposed to promote the learning of generative neural networks, e.g. coarse-to-fine generation [76, 57, 64], feature grids with codebooks [12, 18, 29], tri-planes encoding [69, 13], and primitive embedding [27]. In this paper, similar to [75], we first pre-train a shape point-based AutoEncoder to learn the high-fidelity shape embedding of UDFs, providing a compact latent space for our latent diffusion model to synthesize various realistic shapes.
@@ 3.3 Conditional Shape Generation
Leveraging the cross-attention mechanism, our models capture the mapping between the conditioned input and the geometry implicitly represented by the latent code.
@@ 3.4 Training Details
We trained the Point-based AutoEncoder on Deepfashion3D [79], ShapeNet [6] and Pix3D [62] datasets. The dimension of latent space is $64$. For the input data, we sampled $N=10000$ points on the surface for shape feature extraction. And the UDF values are clipped at $0.1$ for efficiency in both two stages of training. We set $\lambda$ in Eq. 4 to $0.1$. During the diffusion stage, we adopt the DDPM sampling strategy with a sampling step $T=1000$. For the conditioning shape generation, we adopt a frozen CLIP [52] text encoder as text encoder $\tau_{text}$ and employ its image encoder as sketch and image encoder $\tau_{img}$. More details can be found in our Appendix.
@@ 4.3 Conditional Shape Generation
Sketch-guided Shape Generation
[…]
Figure 6: Qualitative results of Sketch Conditioned Generation. Our method generates higher quality and detailed results aligned with input sketch with arbitrary topology.
For sketch-conditioned shape generation, we follow Zheng et al. [76] to generate 2D sketches for each shape in the dataset by rendering its shading images from different views and extracting their edges via the Canny edge detector [3]. The Chamfer distance (CD), Earth mover’s distance (EMD), and the Voxel-IOU are used as evaluation metrics. We compare our
@@
method with two state-of-the-art methods for 3D reconstruction and generation, SDFusion [12] and LAS-Diffusion [76]. Quantitatively, as shown in Tab. 3, our method outperforms other methods on all metrics. As shown in Fig. 6, SDFusion and LAS-Diffusion generate unrealistic results in open surfaces as the drawback of SDF representations, while our method generates high-quality and detailed results aligned with the input sketch.
@@
[…]
4.4 Shape for Simulation
Furthermore, we show that the generated clothing can be easily integrated into a simulation platform [^1] https://www.marvelousdesigner.com/ for virtual try-on, which is shown in Fig. 9. We use the motion sampled from AMASS [43] mocap dataset. Please refer to the video on our webpage in the supplementary material for the visualization. These compelling results validate our high-quality 3D asset generation.
4.5 Grid-based Distance Fields V.S. Continuous Distance Fields
[…]
@@ 4.6 Latent Interpolation
5 Conclusion
In this paper, we present Surf-D for generating high-quality 3D shapes as surfaces with arbitrary topologies based on a diffusion model, leveraging Unsigned Distance Fields (UDF) to represent the surfaces as a general shape representation. For learning and generating detailed geometry, we propose a latent diffusion model for UDFs by first conducting shape embedding to compress various surfaces into a compact latent space and utilizing diffusion models to capture the shape distribution. With our point-based AutoEncoder that allows gradient querying for each query point in a differentiable manner, our method learns detailed geometry at high resolution. Additionally, the curriculum learning is incorporated for efficient learning of diverse surfaces. Surf-D attains superior quality and controllability in shape generation across various modalities. Extensive experiments validate the effectiveness and efficacy of the proposed method.
[…]
Figure 12: Failure case of Surf-D.
Limitations and Future Works Although Surf-D demonstrates superior performance in shape generation across multiple modalities, it has some limitations. As shown in Fig. 12, Surf-D is unable to accurately reconstruct the iron chain (see the zoom-in area). This is because the frequency of geometric details in the iron chain portion is higher than the resolution of the grid for the marching cube, resulting in the loss of geometric details in that part. It would be valuable to explore how to extract mesh from a UDF field efficiently and faithfully. As current mesh extraction methods from the UDF field still require gradients, extracting mesh from the UDF field in a gradient-free way will be an interesting direction. Furthermore, our model requires a relatively large amount of expensive 3D data for capturing the distribution of versatile shapes. Thus, learning shape knowledge from 2D images would significantly alleviate the need for 3D data, which could also be an intriguing research direction.
@@ 5 Conclusion
This supplementary material covers: more experimental results (Sec. 0.A); additional ablation studies (Sec. 0.B); implementation details (Sec.0.C).
@@ 0.B.1 Ablation of Curriculum Learning Strategy
Figure 17: Ablation results of curriculum learning strategy.
[…]
Figure 17: Ablation results of curriculum learning strategy.
Figure 18: Comparison of different reconstruction loss.

### v3

@@ 1 Introduction
Generating a diverse array of 3D shapes using various conditions is an important problem in computer vision and graphics, given its wide applications, such as gaming, filmmaking, and the metaverse.
@@
Recent advancements for 3D shape generation [5, 1, 47, 33, 21, 39, 45, 35] aim to make this process more accessible. The diverse interplay of topology and intricate geometry gives rise to countless 3D shapes in our world. However, existing generation works still suffer from limited topologies and the lack of geometric details. Previous methods present shapes using point clouds [1, 47], voxel representations [68, 88], and signed distance fields (SDFs) [11, 52]. However, point cloud and voxel representations encounter difficulties in faithfully extracting surfaces. Implicit representations, particularly SDFs, excel in describing closed surfaces but face challenges with open surfaces and complex topologies. Despite attempts to enhance them, e.g. 3PSDF [10] and TSDF [71], the representations remain reliant on the underlying definition of SDF. G-Shell [41] is proposed to represent open surfaces but still falters with non-manifold structures and computational constraints.
@@
Moreover, to learn the compact shape latent space efficiently and effectively on diverse surfaces, we employ a curriculum learning strategy [2] to train our model in an easy-to-hard sampling order, i.e. progressively extending the current training dataset during training. This strategy promotes our model to develop a diverse latent space, avoiding entrapment into generating mean shape, i.e. model collapse. As a result, our model is not only lightweight and learning-efficient but also feasible to convert high-quality generated implicit shapes to explicit meshes.
@@
    Surf-D achieves superior performance in generating 3D shapes with diverse topologies while enabling various modalities conditional generation tasks.
@@ 2 Related Work
Tremendous efforts have been made for 3D shape generation [5, 1, 47, 4, 21, 40, 7, 35, 30, 31, 91, 19, 81, 72], given its broad applications, employing various shape representations such as voxel grids [86, 83], point clouds [47, 76, 95], and implicit neural fields [55, 17, 12, 25, 46, 50, 38, 79, 10, 74, 65, 64]. Additionally, various methods have been developed to guide the synthesis with text prompt [63, 18, 43, 9, 42, 78], single-view image [39, 45, 37, 36, 85], sketch [87, 93] and sparse control anchors [48, 13]. In the early stages, voxel-based generation methods [86, 83, 59] can synthesize various coarse shapes. However, the voxel representation can not scale to high-resolution generation [94, 93], due to its cubic space complexity. Utilizing point clouds [47, 76, 95], which are compact and readily available, for 3D object representation can achieve arbitrary shape generation with their well-suitability to learn geometry and texture information [53, 87]. However, it requires complex post-processing to faithfully convert the synthesized point clouds into meshes for real-world applications [57]. The implicit representation such as SDF [12] and differentiable explicit shape representations such as FlexiCubes [65] and DMTet [64] can achieve detailed and high-quality results. However, they perform poorly in modeling the non-watertight shapes as their definitions are based on SDF. Also, they are highly memory-consuming when incorporated with diffusion models.
@@ 2.2 Diffusion Probabilistic Models
Diffusion probabilistic models [24] have demonstrated superior performance in generating samples from a distribution by learning to denoise from a data point gradually [16, 80, 77, 75]. The diffusion models produce high-quality contents including images [39, 16, 54, 80, 61], video [77, 23, 20], and human motion [75, 26, 66, 96]. Previous shape generation works [40, 25] directly train diffusion models on massive grids storing shape representation like SDF, consuming plenty of GPUs and taking days for training. To improve the efficiency, the latent diffusion approach [61] has been explored in 3D shape generation. However, naively applying the latent diffusion model leads to poor generation capability as the vanilla neural field is easy to overfit a single shape. Thus, many efficient generation frameworks are proposed to promote the learning of generative neural networks, e.g. coarse-to-fine generation [93, 67, 76], feature grids with codebooks [12, 18, 32], tri-planes encoding [84, 13], and primitive embedding [28]. In this paper, similar to [92], we first pre-train a shape point-based AutoEncoder to learn the high-fidelity shape embedding of UDFs, providing a compact latent space for our latent diffusion model to synthesize various realistic shapes.
@@ 3.3 Conditional Shape Generation
With the cross-attention, our models capture the mapping between the conditioned input and the geometry implicitly represented by the latent code.
@@ 3.4 Training Details
We trained the Point-based AutoEncoder on Deepfashion3D [97], ShapeNet [6] and Pix3D [73] datasets. The dimension of latent space is $64$. For the input data, we sampled $N=10000$ points on the surface for feature extraction. And the UDF values are clipped at $0.1$ for efficiency in training. We set $\lambda$ in Eq. 4 to $0.1$. During the diffusion stage, we adopt the DDPM sampling strategy with a sampling step $T=1000$. For the conditioning shape generation, we adopt a frozen CLIP [60] text encoder as text encoder $\tau_{text}$ and employ its image encoder as sketch and image encoder $\tau_{img}$. More details are in the Appendix.
@@ 4.3 Conditional Shape Generation
Sketch-guided Shape Generation
[…]
Figure 6: Qualitative results of Sketch Conditioned Generation. Our method generates higher quality and detailed results aligned with input sketch.
For sketch-conditioned shape generation, we follow Zheng et al. [93] to generate 2D sketches for each shape in the dataset by rendering its shading images from different views and extracting their edges via the Canny edge detector [3]. The Chamfer distance (CD), Earth mover’s distance (EMD), and the Voxel-IOU are used as evaluation metrics. We compare our method with two state-of-the-art methods for 3D reconstruction and generation, SDFusion [12] and LAS-Diffusion [93]. Quantitatively, as shown in Tab. 3, our method outperforms other methods on all metrics. As shown in Fig. 6, SDFusion and LAS-Diffusion generate unrealistic results in open surfaces as the drawback of SDF representations, while our method generates high-quality and detailed results aligned with the input sketch.
@@
4.4 Grid-based Distance Fields V.S. Continuous Distance Fields
@@ 4.5 Latent Interpolation
5 Conclusion
In this paper, we present Surf-D for generating high-quality 3D shapes as surfaces with arbitrary topologies based on a diffusion model, leveraging Unsigned Distance Fields (UDF) to represent the surfaces as a general shape representation.
@@ 5 Conclusion
For learning and generating detailed geometry, we propose a latent diffusion model for UDFs by first conducting shape embedding to compress various surfaces into a compact latent space and utilizing diffusion models to capture the shape distribution. With our point-based AutoEncoder that allows gradient querying for each query point in a differentiable manner, our method learns detailed geometry at high resolution. Additionally, the curriculum learning is incorporated for efficient learning of diverse surfaces. Surf-D attains superior quality and controllability in shape generation across various modalities. Extensive experiments validate the effectiveness and efficacy of the proposed method.
Acknowledgements
This work is partly supported by the Innovation and Technology Commission of the HKSAR Government under the ITSP-Platform grant (Ref: ITS/319/21FP) and the InnoHK initiative (TransGP project).
@@ Acknowledgements
This appendix covers: more experimental results (Sec. 0.A); additional ablation studies (Sec. 0.B); implementation details (Sec.0.C); limitation and future work(Sec.0.D);). We highly suggest readers visit the including webpage for a more comprehensive review.
@@ 0.A.3 More Results of Unconditional Shape Generation
0.A.4 Shape for Simulation
Furthermore, we investigate the generated shape for simulation [34, 56]. We show that the generated clothing can be easily integrated into a simulation platform [^1] https://www.marvelousdesigner.com/ for virtual try-on, which is shown in Fig. 15. We use the motion sampled from AMASS [49] mocap dataset. Please refer to the video on our webpage for the visualization. These compelling results validate our high-quality 3D asset generation.
[…]
@@ 0.B.2 Ablation of the Training Loss Function
[…]
Figure 16: Ablation results of curriculum learning strategy.
[…]
Figure 17: Comparison of different reconstruction loss.
@@ Appendix 0.C Implementation Details
Appendix 0.D Limitations and Future Works
[…]
Figure 18: Failure case of Surf-D.
Although Surf-D demonstrates superior performance in shape generation across multiple modalities, it has some limitations. As shown in Fig. 18, Surf-D is unable to accurately reconstruct the iron chain (see the zoom-in area) because the frequency of geometric details in the iron chain portion is higher than the resolution of the grid for the marching cube, resulting in the loss of geometric details in that part. It would be valuable to explore how to extract mesh from a UDF field efficiently and faithfully. As current mesh extraction methods from the UDF field still require gradients, extracting mesh from the UDF field in a gradient-free way will be an interesting direction. Furthermore, our model requires a relatively large amount of expensive 3D data for capturing the distribution of versatile shapes. Thus, learning shape knowledge from 2D images would significantly alleviate the need for 3D data, which could also be an intriguing research direction.